# Uncertainty quantification for intervals

### v1

@@ 1.2 Notation
$G_{n}$ Empirical distribution function of $\Phi_{1},\ldots,\Phi_{n}$
$\psi(\cdot)$ $|\cdot-b|,b\in[0,1]$
$V_{i}$ Unobserved scores, $V_{i}:=\psi(\hat{F}_{1}(T_{i},X_{i}))$
$r_{N-n}$ Uniform rate of convergence for survival curves
2 Related work
2.1 Statistical literature on uncertainty quantification
@@
In recent years, uncertainty quantification became an active research area [25, 55]. The impact of uncertainty quantification on data-driven systems has led to a remarkable surge of interest in both applied and theoretical domains. These works delve into the profound implications of uncertainty quantification in statistics and beyond, such as in the biomedical field [30, 3].
@@
Geisser’s pioneering book [25] develops a mathematical theory of predictive inference. Building upon Geisser’s foundations, Politis presented a comprehensive methodology [55] that effectively harnesses resampling techniques. Additionally, the book of Vovk, Gammmerman and Shafer [69] has been pivotal, see recent reviews [2, 20].
@@
One of the most widely used and robust frameworks for quantifying uncertainty in statistical and machine learning models is conformal inference [59]. The central idea of conformal inference is rooted in the concept of exchangeability [41], even though we will assume that the observedd random elements $\mathcal{D}_{N}$ are independent and identically distributed (i.i.d.). Now, we present a general overview of conformal inference methods for regresion models with scalar responses. Consider the sequence $\mathcal{D}_{N}=\{(X_{i},T_{i})\}^{N}_{i=1}$ of i.i.d. random variables. Given a new i.i.d. random pair $(X,T)$ with respect to $\mathcal{D}_{N}$, conformal prediction provides a family of algorithms for constructing predictive intervals independently of the regression method used, as introduced by [69].
@@
Theorem 2.1 (Full conformal prediction [69]).
@@
The same result holds for split conformal methods, which separates the fitting and ranking steps using sample splitting, and its computational cost is simply that of the fitting step [42].
@@
Conformal inference techniques have been applied to various regression settings, including estimation of the conditional mean [42], conditional quantiles [58], and different functionals of conditional distributions [60, 8]. In recent years, multiple extensions of conformal inference techniques have emerged to handle counterfactual inference problems [7, 74, 36], heterogeneous policy effect [6], reinforcement learning [13], federated learning [48], outlier detection [4], hyphotesis testing [31], robust optimization [37], multilevel structures [19, 15], missing data [51, 77], and survival analysis problems [5, 63], as well as problems involving dependent data such as time series and spatial data [7, 72, 62, 73].
@@
Asymptotic marginal- and in certain cases, conditional- guarantees for the coverage are provided in [79, 71, 11]. The key idea behind these approaches is the application of resampling techniques, originally proposed by [56], to residuals or other score measures. Notably, these methodologies are applicable regardless of the predictive algorithm being used, as emphasized by [55].

### v2

@@ Appendix E Proof of Theorem 3
Let $G_{n}$ be the empirical distribution function of $\Phi_{1},\ldots,\Phi_{n}$ conditional on $\widehat{F}_{1}$ (unknown). Then,
$G_{n}(t)-\mathbb{I}_{n}(t)=-b(t)+o_{P}\left(1\right),\quad t\in[0,1]$
Proof.
$G_{n}(t)-\mathbb{I}_{n}(t)=G_{n}(t)-G^{\prime}(t)+G^{\prime}(t)-\mathbb{I}_{n}(t)=G_{n}(t)-G(t)-b(t)+G^{\prime}(t)-\mathbb{I}_{n}(t)=-b(t)+o_{P}\left(1\right)+o_{P}\left(1\right)=-b(t)+o_{P}\left(1\right)$ because of the classical Glivenko-Cantelli Theorem and weak convergence of the interval process. ∎
We now start te proof of Theorem 3. We denote $P(\cdot)=\mathbb{P}(\cdot\mid\left(X_{i},L_{i},U_{i}\right):i\in\mathcal{I}_{1})$. Denoting by $V_{i}=\psi(\Phi_{i})=\psi(\hat{F}_{1}(T_{i},X_{i}))$ the unobserved scores,
$P\left(T_{N+1}\in\widehat{\mathcal{C}}_{1-\alpha}^{\text{split,*}}\left(X_{N+1}\right)\mid\mathbb{I}_{n}\right)=P\left(\psi(\hat{F}_{1}(T_{N+1},X_{N+1}))\leq\hat{Q}^{*}_{\mathcal{I}_{2}}\mid\mathbb{I}_{n}\right)$
$=P\left(V_{N+1}\leq\hat{Q}^{*}_{\mathcal{I}_{2}}\mid\mathbb{I}_{n}\right)=P\left(\frac{1}{n}\sum_{i=1}^{n}{1}\left\{V^{*}_{i}\geq{V}_{N+1}\right\}>\alpha\mid\mathbb{I}_{n}\right).$
Now for all $v>0$
$V^{*}_{i}\leq v\Longleftrightarrow\left|\Phi^{*}_{i}-\frac{1}{2}\right|<v\Longleftrightarrow\Phi^{*}_{i}<\frac{1}{2}+v\textrm{ and }\Phi^{*}_{i}>\frac{1}{2}-v.$
Therefore,
${1}\left\{V^{*}_{i}\leq v\right\}={1}\left\{\Phi^{*}_{i}<\frac{1}{2}+v\right\}-{1}\left\{\Phi^{*}_{i}<\frac{1}{2}-v\right\}.$
So that
$\frac{1}{n}\sum_{i=1}^{n}{1}\left\{V^{*}_{i}\leq v\right\}=\mathbb{I}^{*}_{n}\left(\frac{1}{2}+v\right)-\mathbb{I}^{*}_{n}\left(\frac{1}{2}-v\right)=:\mathbb{I}^{*}_{n}\left(v^{+}\right)-\mathbb{I}^{*}_{n}\left(v^{-}\right).$
Consequently,
$P\left(v\leq\hat{Q}^{*}_{\mathcal{I}_{2}}\mid\mathbb{I}_{n}\right)=P\left(1+\mathbb{I}^{*}_{n}\left(v^{-}\right)-\mathbb{I}^{*}_{n}\left(v^{+}\right)>\alpha\mid\mathbb{I}_{n}\right).$
If we add $\mathbb{I}_{n}\left(v^{+}\right)$, subtract $1-\mathbb{I}_{n}\left(v^{-}\right)$, and multiply by $-\sqrt{n}$ both sides of the inequality we have for $v>0$
$&P\left(v\leq\hat{Q}^{*}_{\mathcal{I}_{2}}\mid\mathbb{I}_{n}\right)\\ &=P\left(\sqrt{n}\left(\mathbb{I}^{*}_{n}\left(v^{+}\right)-\mathbb{I}_{n}\left(v^{+}\right)\right)-\sqrt{n}\left(\mathbb{I}^{*}_{n}\left(v^{-}\right)-\mathbb{I}_{n}\left(v^{-}\right)\right)<\sqrt{n}\left(1-\alpha+\mathbb{I}_{n}\left(v^{-}\right)-\mathbb{I}_{n}\left(v^{+}\right)\right)\mid\mathbb{I}_{n}\right)\\ &=P\left(\sqrt{n}\left(\mathbb{I}^{*}_{n}\left(v^{+}\right)-\mathbb{I}_{n}\left(v^{+}\right)\right)<\sqrt{n}\left(1-\alpha+\mathbb{I}_{n}\left(v^{-}\right)-\mathbb{I}_{n}\left(v^{+}\right)\right)+\sqrt{n}\left(\mathbb{I}^{*}_{n}\left(v^{-}\right)-\mathbb{I}_{n}\left(v^{-}\right)\right)\mid\mathbb{I}_{n}\right)\\ &\stackrel{{\scriptstyle\text{Cor. 3.}}}{{=}}P\left(\sqrt{n}\left(\mathbb{I}_{n}\left(v^{+}\right)-G^{\prime}\left(v^{+}\right)\right)<\sqrt{n}\left(1-\alpha+\mathbb{I}_{n}\left(v^{-}\right)-\mathbb{I}_{n}\left(v^{+}\right)\right)+\sqrt{n}\left(\mathbb{I}^{*}_{n}\left(v^{-}\right)-\mathbb{I}_{n}\left(v^{-}\right)\right)\mid\mathbb{I}_{n}\right)+o_{P}(1)\\ &=P\left(\sqrt{n}\left(\mathbb{I}_{n}\left(v^{+}\right)-G^{\prime}\left(v^{+}\right)\right)-\sqrt{n}\left(1-\alpha+\mathbb{I}_{n}\left(v^{-}\right)-\mathbb{I}_{n}\left(v^{+}\right)\right)<\sqrt{n}\left(\mathbb{I}^{*}_{n}\left(v^{-}\right)-\mathbb{I}_{n}\left(v^{-}\right)\right)\mid\mathbb{I}_{n}\right)+o_{P}(1)\\ &\stackrel{{\scriptstyle\text{Cor. 3}}}{{=}}P\left(\sqrt{n}\left(\mathbb{I}_{n}\left(v^{+}\right)-G^{\prime}\left(v^{+}\right)\right)-\sqrt{n}\left(1-\alpha+\mathbb{I}_{n}\left(v^{-}\right)-\mathbb{I}_{n}\left(v^{+}\right)\right)<\sqrt{n}\left(\mathbb{I}_{n}\left(v^{-}\right)-G^{\prime}\left(v^{-}\right)\right)\right)+o_{P}(1)\\ &=P\left(\left(\mathbb{I}_{n}\left(v^{+}\right)-G^{\prime}\left(v^{+}\right)\right)-\left(\mathbb{I}_{n}\left(v^{-}\right)-G^{\prime}\left(v^{-}\right)\right)<1-\alpha+\mathbb{I}_{n}\left(v^{-}\right)-\mathbb{I}_{n}\left(v^{+}\right)\right)+o_{P}(1)\\ &\stackrel{{\scriptstyle\text{Lemma 2}}}{{=}}P\left(o_{P}(1)<1-\alpha+\mathbb{I}_{n}\left(v^{-}\right)-\mathbb{I}_{n}\left(v^{+}\right)\right)+o_{P}(1)\\ &=P\left(G_{n}\left(v^{+}\right)-{G}_{n}\left(v^{-}\right)+o_{P}(1)<1-\alpha+\mathbb{I}_{n}\left(v^{-}\right)-{G}_{n}\left(v^{-}\right)-\mathbb{I}_{n}\left(v^{+}\right)+{G}_{n}\left(v^{+}\right)\right)+o_{P}(1)\\ &=P\left(G_{n}\left(v^{+}\right)-{G}_{n}\left(v^{-}\right)<1-\alpha+b(v^{-})-b(v^{+})+o_{P}(1)\right)+o_{P}(1)\\ &=P\left(\frac{1}{n}\sum_{i=1}^{n}{1}\left\{{V}_{i}\leq v\right\}<1-\alpha+\mathcal{O}_{\mathbf{P}}(r_{N-n})+o_{P}(1)\right)+o_{P}(1)\\ \end{aligned}$
Therefore,
$P\left(V_{N+1}\leq\hat{Q}^{*}_{\mathcal{I}_{2}}\mid\mathbb{I}_{n}\right)=$
$=P\left(\frac{1}{n}\sum_{i=1}^{n}{1}\left\{{V}_{i}\leq V_{N+1}\right\}<1-\alpha+\mathcal{O}_{\mathbf{P}}(r_{N-n})+o_{P}(1)\right)+o_{P}(1)$
. The probability in LHS of the previous line involves randomness just due to resampling and the probability in RHS is conditional on the data indexed by $\mathcal{I}_{1}$. Now by taking expectations over the proper training set on both sides of equal sign we arrive to an equality involving $\mathbb{P}$. The rest of the proof goes on as in Theorem 2 in Chernozhukov, Wüthrich & Zhu (2021).
Appendix F Proof of Lemma 1
See the proof for Lemma 1 in Chernozhukov, Wüthrich & Zhu (2021) and use the fact that
[…]
Appendix G Proof of Corollary 4
First,
$\left|{V}_{N+1}-\psi(F_{N+1})\right|=\left|\left|\widehat{F}_{1}(X_{N+1},T_{N+1})-\frac{1}{2}\right|-\left|F(X_{N+1},T_{N+1})-\frac{1}{2}\right|\right|$
$\leq\left|\widehat{F}_{1}(X_{N+1},T_{N+1})-F(X_{N+1},T_{N+1})\right|$
Now,
$\left|\widehat{K}({V}_{N+1})-K(\psi(F_{N+1}))\right|=\left|\widehat{K}({V}_{N+1})-{K}({V}_{N+1})\right|+\left|{K}({V}_{N+1})-K(\psi(F_{N+1}))\right|$
$\leq\mathcal{O}_{P}\left(\operatorname{max}\{r_{N-n}^{2/3},n^{-\frac{1}{2}}\}\right)+W\left|{V}_{N+1}-\psi(F_{N+1})\right|=\mathcal{O}_{P}\left(\operatorname{max}\{r_{N-n}^{2/3},r_{N-n},n^{-1/2}\}\right)$
Then, $\left|\widehat{K}({V}_{N+1})-K(\psi(F_{N+1}))\right|={O}_{P}\left(\operatorname{max}\{r^{2/3}_{N-n},n^{-1/2}\}\right)$ as $r_{N-n}<r_{N-n}^{2/3}$ for $r_{N-n}<1$. Finally, we have
$P\left(1-\widehat{K}({V}_{N+1})>{O}_{P}\left(n^{-\frac{1}{2}}\right)+\alpha\right)+{O}_{P}\left(n^{-\frac{1}{2}}\right)$
$=P\left(1-{K}(\psi(F_{N+1}))>{O}_{P}\left(n^{-\frac{1}{2}}\right)+{O}_{P}\left(\operatorname{max}\{r^{2/3}_{N-n},n^{-1/2}\}\right)+\alpha\right)+{O}_{P}\left(n^{-\frac{1}{2}}\right)$
$=1-\alpha+{O}_{P}\left(\operatorname{max}\{r^{2/3}_{N-n},n^{-1/2}\}\right)$
Appendix H Statistical literature on uncertainty quantification
@@
In recent years, uncertainty quantification became an active research area Geisser (2017), Politis (2015). The impact of uncertainty quantification on data-driven systems has led to a remarkable surge of interest in both applied and theoretical domains. These works delve into the profound implications of uncertainty quantification in statistics and beyond, such as in the biomedical field Banerji et al. (2023).
@@
Geisser’s pioneering book Geisser (2017) develops a mathematical theory of predictive inference. Building upon Geisser’s foundations, Politis presented a comprehensive methodology Politis (2015) that effectively harnesses resampling techniques. Additionally, the book of Vovk, Gammmerman and Shafer Vovk et al. (2005) has been pivotal, see recent reviews Angelopoulos & Bates (2023), Fontana et al. (2023).
@@
One of the most widely used and robust frameworks for quantifying uncertainty in statistical and machine learning models is conformal inference Shafer & Vovk (2008). The central idea of conformal inference is rooted in the concept of exchangeability Kuchibhotla (2020), even though we will assume that the observedd random elements $\mathcal{D}_{N}$ are independent and identically distributed (i.i.d.). Now, we present a general overview of conformal inference methods for regresion models with scalar responses. Consider the sequence $\mathcal{D}_{N}=\{(X_{i},T_{i})\}^{N}_{i=1}$ of i.i.d. random variables. Given a new i.i.d. random pair $(X,T)$ with respect to $\mathcal{D}_{N}$, conformal prediction provides a family of algorithms for constructing predictive intervals independently of the regression method used, as introduced by Vovk et al. (2005).
@@
Theorem 5 (Full conformal prediction Vovk et al. (2005)).
@@
The same result holds for split conformal methods, which separates the fitting and ranking steps using sample splitting, and its computational cost is simply that of the fitting step Lei et al. (2018).
@@
Conformal inference techniques have been applied to various regression settings, including estimation of the conditional mean Lei et al. (2018), conditional quantiles Sesia & Candès (2020), and different functionals of conditional distributions Singh et al. (2023), Chernozhukov, Wüthrich & Zhu (2021). In recent years, multiple extensions of conformal inference techniques have emerged to handle counterfactual inference problems Chernozhukov, Wüthrich & Zhu (2021), Yin et al. (2022), Jin et al. (2023), heterogeneous policy effect Cheng & Yang (2022), reinforcement learning Dietterich & Hostetler (n.d.), federated learning Lu et al. (2023), outlier detection Bates et al. (2023), hyphotesis testing Hu & Lei (2023), robust optimization Johnstone & Cox (2021), multilevel structures Fong & Holmes (2021), Dunn et al. (2022), missing data Matabuena et al. (2022), Zaffran et al. (2023), and survival analysis problems Candès et al. (2023), Teng et al. (2021), as well as problems involving dependent data such as time series and spatial data Chernozhukov, Wüthrich & Zhu (2021), Xu & Xie (2021), Sun (2022), Xu & Xie (2023).
@@
Asymptotic marginal- and in certain cases, conditional- guarantees for the coverage are provided in Zhang & Politis (2022), Wu & Politis (2023), Das et al. (2022). The key idea behind these approaches is the application of resampling techniques, originally proposed by Politis & Romano (1994), to residuals or other score measures. Notably, these methodologies are applicable regardless of the predictive algorithm being used, as emphasized by Politis (2015).